OpenStudy (anonymous):

Write the standard form of the equation of the circle with the given characteristics. Center: (4, −4); solution point: (1, 0)

OpenStudy (anonymous):

what is the radius of the circle?

OpenStudy (anonymous):

Well, i guess (1,0) is on the circle so i guess r=sqrt((1-4)^2+(0+4)^2)

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

(x-4)^2+(y+4)^2=r^2, other people feel free to chime in as to what r is.

OpenStudy (anonymous):

(x-4)^2+(y+4)^2=a^2,where a is the radius

OpenStudy (jamesj):

The general form of the equation of a circle centered at \( (a,b) \) and of radius \(r\) is \[ (x-a)^2 + (x-b)^2 = r^2 \] We have the center of the circle, (a,b) = (4,-4), and now as others are suggesting you need the radius.

OpenStudy (anonymous):

So r=5

OpenStudy (jamesj):

exactly. Hence the equation of the circle is (x-4)^2 + (y-(-4))^2 = 5^2. Now just clean that up a little.

OpenStudy (anonymous):

a=5